OpenStudy (anonymous):

∫√(sec^2x.cosec^2x) dx

myininaya (myininaya):

\[\int\limits\limits_{}^{}\sqrt{\sec^2(x)\csc^2(x)} dx ?\]

myininaya (myininaya):

\[\sqrt{t^2}=t \text{ assuming } t>0\]

myininaya (myininaya):

\[\int\limits_{}^{}\sqrt{(\sec(x) \csc(x))^2} dx\] let me know if you any help after this part

OpenStudy (anonymous):

I need

OpenStudy (freckles):

how far have you gotten on this one?

OpenStudy (freckles):

here is a hint you might find handy sin(2x)=2sin(x)cos(x) or 1/2*sin(2x)=sin(x)cos(x)